OpenStudy (anonymous):

Factor the polynomial: 42w10 + 24w6 PLEEAASSEE HELP !

OpenStudy (anonymous):

What is your greatest common factor?

OpenStudy (anonymous):

Of the 42 and 24?

OpenStudy (anonymous):

Yes and then of the w^10 and w^6

OpenStudy (anonymous):

The GCF is 3 right? And Then The GCF for the variables is 3 right? So The Answer Is 3w?

OpenStudy (anonymous):

Here is how I like to do it. 42 = 2 * 21 42 = 2 * 3 * 7 24 = 2 * 12 24 = 2 * 2 * 6 24 = 2 * 2 * 2 * 3 so 42 = 2 * 3 * 7 24 = 2 * 2 * 2 * 3 they both have a 2 and a 3 so the GCF = 6 When you have variables and exponents, you take the lowest exponent so w^6 SO the GCF is 6w^6

OpenStudy (anonymous):

But Theses Are The Multiple Choice? So If The Answer Is 6w^6, It's Not Here. . . =( A 2x(x2 + 2x + 4) C 2x(x2 + 2x + 8) B 2x(x + 2)(x + 4) D 2x3 + 4x2 + 8x